Express the following as cm using decimals.
83 mm ?????????

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

83 mm = 83 / 10 cm
= 8.3 cm

Thanks
Shraddha Singh
3 years ago
Dear student,
 
We know that,
1 mm = 1/10 cm
 
83 mm = 83 × 1/10 cm
 
             = 83/10 cm
 
              = 8.3 cm
 
Hope it helps!!
Thank you ✌️